OpenStudy (anonymous):

Classify the triangle by its angles based on the side lengths 4.2, 6.4, and 7.6

OpenStudy (anonymous):

Acute Right Scalene Obtuse

OpenStudy (anonymous):

scalen since the sides are not equal and we dont know the angles to determine if it were obtuse or acute

OpenStudy (anonymous):

Let "c" be the longest side on each set of three numbers. If c^2 = a^2+b^2, the triangle is right If c^2 > a^2 + b^2, the triangle is obtuse If c^2 < a^2 + b^2, the triangle is acute.

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

So what is the longest side?

OpenStudy (anonymous):

It's 7.6, right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

What is 4.2^2?

OpenStudy (anonymous):

What is 6.4^2?

OpenStudy (anonymous):

What is 7.6^2?

OpenStudy (anonymous):

my bad i read the question wrong lol

OpenStudy (anonymous):

17.6, 40.9, and 57.7

OpenStudy (anonymous):

It's good haha

OpenStudy (anonymous):

a=4.2^2=17.64 b=6.4^2=40.96 c=7.6^2=57.76 Add A and B.

OpenStudy (anonymous):

After you find the sum of A and B, compare it to C.

OpenStudy (anonymous):

58.6

OpenStudy (anonymous):

Is 58.6 equal to C, which is 57.76?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Then what is it? Is is more than C or less than C?

OpenStudy (anonymous):

wouldnt u use sohcahtoa

OpenStudy (anonymous):

to find the angles

OpenStudy (anonymous):

Just what you said @joquez14 , you only use sohcahtoa to find the angles. Look up the question again. What is the given?

OpenStudy (anonymous):

Side lengths, right? So you use the Pythagorean theorem. This is like the easiest lesson in Geometry.

OpenStudy (anonymous):

this question has me totally confused im thinking u had to find the angles to classify it

OpenStudy (sepeario):

another way would've been to use the law of sines to work out the angles.

OpenStudy (anonymous):

@NICK702222 you still there?

OpenStudy (anonymous):

Yeah

OpenStudy (sepeario):

so @joquez14 if you want to evaluate the angles you can also do that.

OpenStudy (anonymous):

I'm super confused

OpenStudy (anonymous):

Just answer my question. :)

OpenStudy (anonymous):

Is the sum of A and B greater than, less than, or equal to C?

OpenStudy (anonymous):

is sine not apart of sohcahtoa

OpenStudy (anonymous):

@Sepeario Just trying to help her find the easiest solution. :)

OpenStudy (anonymous):

Sine, in fact, is part of the SohCahToa. S stands for Sine.

OpenStudy (anonymous):

thats wat i thought so i wasnt completly wrong then

OpenStudy (anonymous):

Is 58.6 equal to 57.7? Is 58.6 more than 57.7? Is 58.6 less than 57.7? @NICK702222

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Obtuse?

OpenStudy (anonymous):

lol nope thts hilarious

OpenStudy (anonymous):

No no wait

OpenStudy (anonymous):

If c=sum, the triangle is right If c>sum, the triangle is obtuse If c<sum, the triangle is acute.

OpenStudy (anonymous):

You're almost there. :)

OpenStudy (anonymous):

First use the distance formula to find the sides. After using the distance formula, you can use Cosine Law to find the angles. But for this problem, one just needs to use the distance formula to find the sides and then graph it and they they tell what type of triangle and to check if it is a 90 or not they can use \( a^2+b^2=^c \) but one can also use this to tell for the other too, but you might be able just to eye it to tell.

OpenStudy (anonymous):

If 57.7=58.6, the triangle is right If 57.7>58.6, the triangle is obtuse If 57.7<58.6, the triangle is acute.

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

So it's acute

OpenStudy (anonymous):

Got it right! I'd really appreciate it if you would give me a medal. :)

OpenStudy (anonymous):

Thanks for the help

OpenStudy (anonymous):

You're welcome!